# Towards Autonomous Driving with Small-Scale Cars: A Survey of Recent Development

### v1

@@ V Future trends
Firstly, it is imperative to contemplate transforming the small-scale platform into a more widely accessible resource for educational purposes and academic research. This necessitates the establishment of a lower entry barrier, reduced pricing, and the implementation of a comprehensive learning pipeline. For educational usage, the first consideration should be given to the ease of learning and maintenance for teachers. A more user-friendly starting point will likely foster increased enthusiasm among educators, encouraging them to incorporate the platform into their daily teaching activities. This, in turn, will provide students with a challenging learning experience, ultimately preparing them for the intricacies of fully autonomous driving in the near future. For research applications, an in-depth exploration of the accessibility of the small-scale car platform is essential. This entails tailoring entry levels to accommodate various research goals, catering to individuals with zero experience in robotics to seasoned professionals. The platform should not only serve as a testbed for autonomous systems but also ignite enthusiasm for exploring diverse robotic configurations. While some platforms currently offer support for different entry levels, a more comprehensive development is warranted to meet the diverse needs of the research community.
@@
Secondly, regarding the more academic research consumption, the versatility of the platforms should be more considered. Specifically, concerning individual small-scale cars, attention should be directed towards enhancing their functional dimensions. In line with advancements in semiconductor technology, where computation and sensor units are becoming more compact yet powerful, platforms should incorporate more sophisticated sensors, thereby augmenting the overall capabilities of the system. Expanding the scope of research considerations, we propose a thorough examination of smart city configurations. While smart cities hold significant potential in advancing research on autonomous driving for small-scale cars, there exists a gap in terms of accessibility, reproducibility, and standardization of best practices. There is an urgent need for a common framework across the research community, encompassing both hardware and software aspects. Furthermore, existing smart city setups often overlook key elements such as weather conditions and pedestrian interactions, which are important in real-world driving scenarios. Addressing these factors in future research is essential for developing comprehensive and realistic autonomous driving solutions within smart city environments.
@@
Thirdly, achieving fully autonomous driving necessitates the integration of critical technologies such as Vehicle-to-Vehicle (V2V), Vehicle-to-Infrastructure (V2I), and Vehicle-to-Everything (V2X). While these technologies have been extensively researched in the literature, there remains a noticeable gap in discussions of small-scale car platforms. Consequently, the community needs to pivot towards exploring and advancing V2V and V2I communications as the next steps in the pursuit of comprehensive autonomous driving solutions.
@@ VI Conclusion
In this paper, we offer an overview of the current state-of-the-art developments in small-scale autonomous cars. Through a detailed exploration of both past and ongoing research in this domain, we illuminate the promising trajectory for the advancement of autonomous driving technology with small-scale cars. We initially enumerate the presently predominant small-scale car platforms widely employed in academic and educational domains and present the configuration specifics of each platform. Similar to their full-size counterparts, the deployment of hyper-realistic simulation environments is imperative for training, validating, and testing autonomous systems before real-world implementation. To this end, we show the commonly employed universal simulators and platform-specific simulators. Furthermore, we provide a detailed summary and categorization of tasks accomplished by small-scale cars, encompassing localization and mapping, path planning and following, lane-keeping, car following, overtaking, racing, obstacle avoidance, and more. Within each benchmarked task, we classify the literature into distinct categories: end-to-end systems versus modular systems and traditional methods versus ML-based methods. This classification facilitates a nuanced understanding of the diverse approaches adopted in the field. The collective achievements of small-scale cars are thus showcased through this systematic categorization. Since this paper aims to provide a holistic review and guide, we also outline the commonly utilized in various well-known platforms. This information serves as a valuable resource, enabling readers to leverage our survey as a guide for constructing their own platforms or making informed decisions when considering commercial options within the community.

### v2

@@ V-A Enhancing Accessibility and Usability
It is imperative to contemplate transforming the small-scale platform into a more widely accessible resource for educational purposes and academic research. This necessitates the establishment of a lower entry barrier, reduced pricing, and the implementation of a comprehensive learning pipeline. For educational usage, the first consideration should be given to the ease of learning and maintenance for teachers. A more user-friendly starting point will likely foster increased enthusiasm among educators, encouraging them to incorporate the platform into their daily teaching activities. This, in turn, will provide students with a challenging learning experience, ultimately preparing them for the intricacies of fully autonomous driving in the near future. For research applications, an in-depth exploration of the accessibility of the small-scale car platform is essential. This entails tailoring entry levels to accommodate various research goals, catering to individuals with zero experience in robotics to seasoned professionals. The platform should not only serve as a testbed for autonomous systems but also ignite enthusiasm for exploring diverse robotic configurations. While some platforms currently offer support for different entry levels, a more comprehensive development is warranted to meet the diverse needs of the research community.
V-B Improving Versatility and Advanced Technology Adoption
Regarding the more academic research consumption, the versatility of the platforms should be more considered. Specifically, concerning individual small-scale cars, attention should be directed towards enhancing their functional dimensions. In line with advancements in semiconductor technology, where computation and sensor units are becoming more compact yet powerful, platforms should incorporate more sophisticated sensors, thereby augmenting the overall capabilities of the system. As discussed in Section IV, most of the current techniques applied in small-scale car platforms are old methods developed for full-scale cars, with simplified apply conditions. With the rapid advancement of AD research, increasingly sophisticated techniques in sensing, perception, and control are being developed for full-scale vehicles. These novel methods typically require advanced sensor suites, including 3D LiDAR and cameras, which demand significant computational resources and power supply—requirements that most small-scale car platforms cannot meet. Additionally, deploying these techniques in real-world systems necessitates fast communication between components. As a result, more powerful platforms are becoming increasingly attractive.
V-C Bridging Gaps in Smart City Configurations
@@
We propose a thorough examination of smart city configurations. While smart cities hold significant potential in advancing research on autonomous driving for small-scale cars, there exists a gap in terms of accessibility, reproducibility, and standardization of best practices. There is an urgent need for a common framework across the research community, encompassing both hardware and software aspects. By addressing these two aspects, small-scale car platforms can more closely resemble real-world AD systems. It unlocks the potential for deploying widely used open-source AD systems, such as Apollo and Autoware, on small-scale car platforms. These AD systems necessitate a comprehensive ecosystem encompassing sensing, perception, localization, planning, and control. Furthermore, existing smart city setups often overlook key elements such as weather conditions and pedestrian interactions, which are important in real-world driving scenarios. Addressing these factors in future research is essential for developing comprehensive and realistic autonomous driving solutions within smart city environments.
V-D Advancing V2V and V2I Integration
@@
Achieving fully autonomous driving necessitates the integration of critical technologies such as V2V, V2I, and V2X. While these technologies have been extensively researched in the literature, there remains a noticeable gap in discussions of small-scale car platforms. Consequently, the community needs to pivot towards exploring and advancing V2V and V2I communications as the next steps in the pursuit of comprehensive autonomous driving solutions.
@@ VI Conclusion
In this survey, we offer an overview of the current state-of-the-art developments in small-scale autonomous cars. Through an in-depth exploration of past and ongoing research, we identify critical challenges and highlight the promising trajectory for advancing small-scale autonomous driving technology. We begin by enumerating the presently predominant small-scale car platforms widely employed in academic and educational domains, detailing the configurations and specifications of each. Similar to their full-scale counterparts, the deployment of hyper-realistic simulation environments is imperative for training, validating, and testing autonomous systems before real-world implementation. To this end, we show the commonly employed universal simulators and platform-specific simulators. Furthermore, we provide a detailed summary and classify the literature into distinct categories: end-to-end systems versus modular systems and traditional methods versus ML-based methods. This classification facilitates a nuanced understanding of the diverse approaches adopted in the field. We introduce methods used for perception, path planning, control, and end-to-end driving. To provide a holistic guide for researchers and practitioners, we also outline the commonly utilized components and tools across various well-known platforms. This information serves as a valuable resource, enabling readers to leverage our survey as a guide for constructing their own platforms or making informed decisions when considering commercial options within the community.